OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 4 x - 2y = -5 2 -2 1 -1

OpenStudy (foreverandalways13):

do you know how to do substitution???

OpenStudy (foreverandalways13):

i spelled that wrong:D

OpenStudy (anonymous):

no all I know is ellimination

OpenStudy (foreverandalways13):

oh thats fine i know that way too:)

OpenStudy (anonymous):

you think you could help me(:

OpenStudy (foreverandalways13):

yup!

OpenStudy (foreverandalways13):

ummmm i think what you have to do is plug like each one in for y and solve for x...i think :D

OpenStudy (foreverandalways13):

@jdoe0001

OpenStudy (foreverandalways13):

he can help hopefully:)

OpenStudy (anonymous):

okay thank you :)

OpenStudy (foreverandalways13):

np :D

OpenStudy (anonymous):

@jaderbrown?

OpenStudy (jdoe0001):

$$ 2x + 3y = 4\\ x - 2\color{red}{y} = -5 \\ \text{solve the 2nd for "y"}\\ y=\color{red}{\frac{x+5}{2}}\\ \text{now that we know what "y" equals to}\\ \text{let us use that and substitute in the 1st one}\\ 2x + 3y = 4\\ x - 2\pmatrix{\color{red}{\frac{x+5}{2}}} = -5 \\ \text{now just solve the 2nd one for "x"} $$

OpenStudy (anonymous):

Cool xD

OpenStudy (jdoe0001):

wait, I substitute it to the wrong one :/

OpenStudy (anonymous):

Hahaha :D

OpenStudy (foreverandalways13):

thanks!!!

OpenStudy (jdoe0001):

$$ 2x + 3y = 4\\ x - 2\color{red}{y} = -5 \\ \text{solve the 2nd for "y"}\\ y=\color{red}{\frac{x+5}{2}}\\ \text{now that we know what "y" equals to}\\ \text{let us use that and substitute in the 1st one}\\ 2x + 3\pmatrix{\color{red}{\frac{x+5}{2}}} = 4\\ x - 2y = -5\\ \text{now just solve the 2nd one for "x"} $$

OpenStudy (jdoe0001):

there

OpenStudy (anonymous):

eliminate y for either equation

OpenStudy (anonymous):

sorry I went to the store, so I got x=2y-5